OpenStudy (chaise):

Limit question. lim t->0 of (e^t+t-1)/t^2 Is this a question that cannot be evaluated with limit tricks, if so, how?

OpenStudy (chaise):

\[\lim_{t \rightarrow 0} \frac{e^t+t-1}{t^2}\]

OpenStudy (anonymous):

There's l'Hospital's rule... does that count as a limit trick?

OpenStudy (chaise):

I would get: \[\lim_{t \rightarrow 0} \frac{e^t+1}{2t}\] When evaluated you get 2/0 Should I try do L'Hopital's rule again?

OpenStudy (anonymous):

You can't do l'Hopital's again because it's not indeterminate form.

OpenStudy (chaise):

That is what I was thinking

OpenStudy (anonymous):

I would check if it actually exists at this point.

OpenStudy (chaise):

I know the answer is infinity. I just don't know how/why

OpenStudy (anonymous):

You sure about that?

OpenStudy (anonymous):

For small negative values it tends towards \(-\infty\) doesn't it?

OpenStudy (chaise):

I hope so, if not that means my book is wrong, and also wolframalpha.

OpenStudy (anonymous):

Yeah, when you have an asymptote that means the limit doesn't exist bro.

OpenStudy (chaise):

Cool. So how do I know there is an asymtope just by looking at the equation?

OpenStudy (anonymous):

\[ \lim_{t\to 0^+}f(t) = \infty \neq -\infty =\lim_{t\to 0^-}f(t) \]

OpenStudy (anonymous):

When you suspect a limit doesn't exist, your only choice is really to plug in really close values.

OpenStudy (anonymous):

For example.... plug in 0+(000.1) and 0-(000.1)

OpenStudy (chaise):

Ah, okay. Thanks I guess. ^_^

OpenStudy (anonymous):

Really, you underestimate how magical limits are...

OpenStudy (anonymous):

There is more magic in limits then in all of Equestria. If you could do a limit as trivially as you could do multiplication, most of calculus would be a joke.

OpenStudy (chaise):

Haha, I guess so. Thanks for your help.